# Continual Learning in Medical Imaging Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v1

@@ 2.4 Continual learning pipeline
The design of a CL pipeline is illustrated in  Figure 1 which provides an overview of the key components and stages involved in the construction of the CL pipeline. Given a problem statement, first, we need to identify which CL scenarios it falls under, i.e., whether there is a possibility of domain shifts in future data, inclusion of new classes, or the end application, i.e., the task itself may change. For example, if we want to develop a breast cancer classification model to be able to work on the H&E dataset from different centers then the datasets across centers may have drift and hence fall into the domain incremental scenario of CL. Detailed information about CL scenarios is provided in Section 3. Once we have identified the CL scenario, training and testing datasets need to be prepared to mimic a continual stream of datasets arriving one after another. The sequence of datasets is frequently referred to as tasks or episodes. Since the term ‘task’ is also used for end application and thus may be challenging to get the true interpretation of the term, hence, we use the ‘episode’ term to refer to the sequence of datasets in a CL pipeline. For each episode, separate training and testing data is prepared, thus to develop and evaluate a CL model for four sequential datasets, the pipeline requires four train-test pairs. Once the datasets are ready, a CL strategy suitable to the task at hand is identified and deployed in any off-the-shelf deep classification or segmentation model. There are various CL strategies, some offering privacy-preserved learning, while some offer better performance but at the cost of more resources, storage, and privacy violations as they store some past data. Typically, the model is trained on the first training episode and evaluated on testing data from all the episodes. After this the training shifts to the next episode where the inclusion of partial training data from the previous episode is possible. Here, after training, the evaluation is again done on all the testing data and the process repeats until the last episode. Application-specific performance metrics (e.g., accuracy, dice similarity coefficient, etc.) on testing data of each episode are observed and analyzed over the sequence. Then, we can compute various metrics on top of it to quantify forgetting and forward transfer. Lastly, the CL framework is evaluated against state-of-the-art (SOTA), algorithms offering upper and lower bounds of performance. Joint or cumulative training gives the highest average performance whereas naively finetuning on the current episode gives the lowest performance .
@@ 2.5 Application of continual learning
Continual learning, also known as incremental learning, has numerous applications across various domains due to its ability to mitigate catastrophic forgetting. The efficiency and robustness of CL in real-time scenarios derive from its ability to adapt, reduce computational overhead, and address the challenges of dynamic data. This makes it a valuable approach for applications where time constraints, adaptability, and efficiency are of great importance. In the medical domain, CL has been widely explored for various segmentation and classification applications (Figure 2) and continuously exhibited its merits over static models as reflected by the increasing number of research contributions (Figure 3). CL can improve the diagnosis and decision-making ability in a resource-constrained environment. Further, it is beneficial in real-time monitoring of patients, telemedicine, and maintaining a dynamic knowledge base. This section presents the most prominent applications of continual learning in medical settings in terms of the number of publications.
@@
Radiology and imaging: Radiological imaging techniques are continually advancing, requiring models to adapt to new technologies and methodologies. For CT, research has investigated the impact of scanners and reconstruction parameters on both machine learning predictions and human annotations. The findings indicate that the variability introduced by different scanners has a detrimental effect on radiomics  Mackin et al. (2015)  Prayer et al. (2021) and other imaging features. Therefore, it is imperative to take into account this scanner variability when developing machine learning models. Continual learning facilitates the dynamic adjustment of models to evolving imaging techniques, ensuring accurate interpretations  Li and Hoiem (2018)  Zhang et al. (2023a). Table 13 and Table 14 present the selected works for continual learning based classification and segmentation in the medical domain.
@@
Disease progression modeling: Continual learning is crucial in tracking the progression of diseases such as Alzheimer’s Disease (AD). Models need to adapt to new patient data over time, incorporating the latest diagnostic criteria and treatment strategies Li et al. (2020). One research proposed a method for AD progression in a continual learning manner that respects the longitudinal data sets coming in sequence and ensures equal prediction accuracy for future visits  Zhang and Wang (2019).
@@
Drug discovery: Continual learning can accelerate drug discovery by continuously integrating newly discovered chemical compounds, pharmacological data, and clinical trial results to predict drug interactions and efficacy. A recent work presented a continual learning-based model called Multi-Scale Temporal Convolutional Networks-based AntiBacterial Peptide prediction (MSTCN-ABPpred). This model is designed for the classification and discovery of antibacterial peptides (ABPs). While MSTCN-ABPpred can dynamically adapt and retrain based on predicted ABPs and non-ABPs within protein sequences, it does have limitations in terms of providing information about the identified ABPs’ targets, haemotoxicity, cytotoxicity, and minimum inhibitory concentration (MIC) against various bacteria  Singh et al. (2023).
@@ 3 Continual learning scenarios
Depending upon what kind of change in the sequence of data is expected, the existing CL scenarios can be broadly categorized into five categories, viz., (a) data-incremental, (b) class-incremental, (c) task-incremental, (d) domain-incremental, and (e) hybrid. Figure 4 shows the percentages of the above-mentioned incremental CL scenarios exploited in recent years. The statistics reflect that out of four major settings, domain-incremental is highly explored with a 51.9% ratio followed by class-incremental settings having 26.9% works and data-incremental being the least challenging and hence rarely explored for different medical image analysis applications.
Table 2: List of various data incremental scenarios in literature
@@
Ravishankar et al. (2019)
Pneumothorax identification from X-ray images (2 Ep.)
4 batches each of 2K were created from a subset of ChestXRay (Wang et al. 2017)
Kaustaban et al. (2022)
Histopathology data-based tumor classification (4 Ep.)
Ep. was created in data incremental fashion from colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes, (i.e., all 9 classes are present in each episode)
3.1 Data incremental scenarios
Data incremental scenario involves the continuous learning process where the model keeps adapting to incoming data streams that come from the same source. This scenario doesn’t involve dealing with entirely new categories or significant changes in data patterns, making it generally the least challenging among all other continual learning scenarios. In the medical domain, especially digital pathology, a common practice is for expert pathologists to annotate a dataset in multiple stages or batches Kaustaban et al. (2022). Each of these annotation batches can be thought of as representing samples from the same underlying data distribution, ensuring some consistency in the data source across different batches. Thus, data incremental scenarios deal with the condition where the samples of a dataset are sequentially made available. Owning to its simplicity, this category is not very popular. Only a few classification applications are explored in this scenario, as mentioned in Table 2. For example, in the breast cancer classification application, Chen and Tang (2022) set up data incremental scenarios on the BreakHis Spanhol et al. (2015) dataset. More and more samples from each of the malignant and benign classes are added over the subsequent episodes.  Ravishankar et al. (2019) try to exhibit a scenario for pneumothorax detection in a hospital where data arrives in an incremental fashion. They sequentially fed 2000 samples in 4 subsequent episodes from an X-ray dataset for Pneumothorax identification consisting of a total of 8000 samples.
@@ 3.1 Data incremental scenarios
Incrementally adding anatomical structure, private data
Ozdemir and Goksel (2019)
MRI based knee segmentation (2 Ep.)
@@
SKI10 MICCAI Grand Challenge (Heimann et al. 2010)
Li et al. (2020)
Dermoscopic images based skin disease classification (4-20 Ep.)
@@
(a) CIS for Skin8 dataset having total 8 classes on which 2 classes per episode used, (b) CIS for Skin40 dataset from ISIC2019 (Tschandl et al. 2018) consists of 40 classes of skin disease images collected from the internet, here 2/5/10 classes per episode used in 3 exp, (c) non-medical dataset CIFAR100 with 100 classes, here 5/10/20 classes per episode used in 3 exp
Liu et al. (2022)
Class incremental segmentation of abdomen organs (4 Ep.)
@@
4 Ep. created for incremental segmentation of liver, spleen, pancreas, right kidney, and left kidney organs from CT datasets (Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
Derakhshani et al. (2022)
CIS: disease classification (4 Ep.)
@@
4 Ep. created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (taken from MedMNIST repository (Yang et al. 2023a))
Akundi and Sivaswamy (2022)
Chest X-ray classification in CIS (5 Ep.)
@@
5 classes from CheXpert dataset
Kaustaban et al. (2022)
Histopathology data based tumor classification (4 Ep.)
@@
4 Ep. were created in CIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total 9 classes
Chen and Tang (2022)
@@
Histopathology data-based breast cancer classification (4 Ep.)
4 Ep. created in CIS from BreakHis dataset
Zhang et al. (2023b)
(a) CT based abdomen, gastrointestinal, and other organ segmentation (7, 3, 4 classes in 3 Ep.), (b) abdomen segmentation to liver tumor segmentation task (13 classes, 1 class in 2 Ep.)
@@
(a) Abdomen segmentation from JHH (Xia et al. 2022) (private) (b) 13 class abdomen segmentation dataset as BTCV (Landman et al. 2015) and liver tumor segmentation data as LiTS (Bilic et al. 2023)
Ji et al. (2023)
3D CT scan based whole-body organs segmentation (4 Ep. in 2 permutations of episodes)
@@
TotalSegmentator (Wasserthal et al. 2023) (103 classes: various organs), 3 private datasets: ChestOrgan (31 classes: chest scans), HNOrgan (13 classes: head and neck scans), EsoOrgan (1 class: esophageal cancer)
Chee et al. (2023)
Cancer classification (3,5 Ep.), diabetic retinopathy classification (2 Ep.), skin lesions classification (3,4 Ep.)
@@
(a) Histopathology data based human colorectal cancer classification: CCH5000 (Kather et al. 2016) (total 8 classes, Exp=(4,1,1,1,1), Exp=(4,2,2)), (b) diabetic retinopathy (DR) classification using retinal images: EyePACS (kag ) (total 5 classes: no DR, mild DR, moderate DR, severe DR and proliferative DR, Exp=(3,2)), (c) pigmented skin lesions classification: HAM10000 (Tschandl et al. 2018) (total 7 classes: Exp=(4,1,1,1), Exp=(4,2,1))
Zhang et al. (2023c)
Disease classification in CIS (4-10 Ep.)
@@
(a) skin lesions classification with Skin8 dataset (International Skin Imaging Collaboration (ISIC) (Tschandl et al. 2018)) having 8 classes distributed over 4 Ep., (b) Path16, a pathology image collection from various public sources: total 16 classes distributed over 7 episodes, (c) CIFAR100: non-medical images, 100 classes distributed over 5-10 episodes
(Bai et al. 2023)
Surgical visual-question localized-answering
EndoVis18, EndoVis17, M2CAI
Sadafi et al. (2023)
Microscopic images based WBC classification (4 Ep.)
@@
With 3 different datasets, 3 separate class-incremental scenarios each having 4 episodes: (a) (3,3,3,4) classes from Matek-19 (Matek et al. 2019), (b) (3,3,3,4) classes from INT-20, and (c) (3,2,2,3) classes from Acevedo-20 (Acevedo et al. 2020)
@@ 3.2 Class incremental scenarios
Class Incremental Scenario (CIS) refers to the situation where we aim to adapt the model to accommodate novel classes from new data streams. Given that each batch of these unseen classes offers knowledge that largely differs from the previous batches, this scenario is anticipated to be the most challenging among continual learning scenarios. An instance of this challenge is incorporating new types of tissues into a tumor classification model or introducing novel cell types into a model designed for AI-assisted diagnosis in cell detection. In this scenario, there is only one incremental task that gets new additions of classes in subsequent episodes. Some assume mutually exclusive classes whereas some overlapping classes.
Mainly multi-class datasets with more than 5-6 classes are considered to create a class incremental scenario. There can be an equal or variable number of classes across the episodes. Chen and Tang (2022) propose a breast cancer classification model where they iteratively include malignant and benign classes with one sub-class from each category. This scenario is very popular for segmentation applications where the model may get demand to segment more and more classes over time. An exhaustive list of class incremental scenarios explored in literature is provided via Table 3. References along with detailed segmentation or classification applications are tabulated.
@@
Baweja et al. (2018)
MRI based normal brain structures segmentation (2 Ep.)
@@
Ravishankar et al. (2019)
Chest x-ray view classification (2 Ep.)
Shift from task (4ch vs. PLAX) to task (2ch vs. PSAX)
Kaustaban et al. (2022)
Histopathology data based tumor classification (4 Ep.)
4 Ep. were created in TIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes
Derakhshani et al. (2022)
Disease classification (4 Ep.)
4 Ep. created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (MedMNIST repository (Yang et al. 2023a))
Bera et al. (2023)
Three tasks involving different organs for three different binary segmentation tasks (3 Ep.)
MRI data for binary prostate, spleen, hippocampus segmentation: Promise12 (Litjens et al. 2014) (prostate) $\rightarrow$ MSD (NCI ) (spleen) $\rightarrow$ Drayd (Denovellis et al. 2021) (hippocampus)

### v3

@@ 2.4 Continual learning pipeline
The design of a CL pipeline is illustrated in  Figure 1 which provides an overview of the key components and stages involved in the construction of the CL pipeline. Given a problem statement, first, we need to identify which CL scenarios it falls under, i.e., whether there is a possibility of domain shifts in future data, inclusion of new classes, or the end application, i.e., the task itself may change. For example, if we want to develop a breast cancer classification model to be able to work on the H&E dataset from different centers then the datasets across centers may have drift and hence fall into the domain incremental scenario of CL. Detailed information about CL scenarios is provided in Section 3. Once we have identified the CL scenario, training, and testing, datasets need to be prepared to mimic a continual stream of datasets arriving one after another. The sequence of datasets is frequently referred to as tasks or episodes in literature. We also use these two terms interchangeably in this manuscript. For each episode, separate training and testing data is prepared; thus, the pipeline requires four train-test pairs to develop and evaluate a CL model for four sequential datasets. Once the datasets are ready, a CL strategy suitable to the task at hand is identified and deployed in any off-the-shelf deep classification or segmentation model. There are various CL strategies, some offering privacy-preserved learning, while some offer better performance but at the cost of more resources, storage, and privacy violations as they store some past data. Typically, the model is trained on the first training episode and evaluated on testing data from all the episodes. After this, the training shifts to the next episode, where the inclusion of partial training data from the previous episode is possible. Here, updating the model with new training data, the evaluation is again done on all the testing data, and the process repeats until the last episode. Application-specific performance metrics (e.g., accuracy, dice similarity coefficient, etc.) computed on testing data of each episode are observed and analyzed over the sequence. Then, we can compute various metrics on top of it to quantify forgetting and forward transfer. Lastly, the CL framework is evaluated against state-of-the-art works and non-CL methods offering upper and lower bounds of performance. Joint or cumulative training gives the highest average performance, whereas naively finetuning on the current episode gives the lowest performance Kaustaban et al. (2022); Lenga et al. (2020).
@@ 2.5 Application of continual learning
CL has numerous applications across various domains due to its ability to mitigate catastrophic forgetting. The efficiency and robustness of CL in real-time scenarios derive from its ability to adapt, reduce computational overhead, and address the challenges of dynamic data. This makes it a valuable approach for applications where time constraints, adaptability, and efficiency are of great importance. In the medical domain, CL has been widely explored for various segmentation and classification applications (Figure 2) and continuously exhibited its merits over static models as reflected by the increasing number of research contributions over the year (Figure 3). CL can improve the diagnosis and decision-making ability in a resource-constrained environment. Further, it is beneficial in real-time monitoring of patients, telemedicine, and maintaining a dynamic knowledge base. This section presents the most prominent applications of CL in medical settings in terms of the number of publications.
@@
Radiology and imaging: Radiological imaging techniques are continually advancing, requiring models to adapt to new technologies and methodologies. For CT, research has investigated the impact of scanners and reconstruction parameters on both machine learning predictions and human annotations. The findings indicate that the variability introduced by different scanners has a detrimental effect on radiomics Mackin et al. (2015); Prayer et al. (2021) and other imaging features. Therefore, it is imperative to take into account this scanner variability when developing machine learning models. CL facilitates the dynamic adjustment of models to evolving imaging techniques, ensuring accurate interpretations  Li and Hoiem (2018)  Zhang et al. (2023a). Table 13 and Table 14 present the selected works for CL-based classification and segmentation in the medical domain.
@@
Disease progression modeling: CL is crucial in tracking the progression of diseases such as Alzheimer’s Disease (AD). Models must adapt to new patient data over time, incorporating the latest diagnostic criteria and treatment strategies Li et al. (2020). Recent work was proposed for modeling AD progression in a CL manner that respects the longitudinal data sets coming in sequence and ensures equal prediction accuracy for future visits Zhang and Wang (2019).
@@
Drug discovery: CL can accelerate drug discovery by continuously integrating newly discovered chemical compounds, pharmacological data, and clinical trial results to predict drug interactions and efficacy. A recent work presented a CL-based model called Multi-Scale Temporal Convolutional Networks-based AntiBacterial Peptide prediction (MSTCN-ABPpred). This model is designed for the classification and discovery of antibacterial peptides (ABPs). While MSTCN-ABPpred can dynamically adapt and retrain based on predicted ABPs and non-ABPs within protein sequences, it does have limitations in terms of providing information about the identified ABPs’ targets, haemotoxicity, cytotoxicity, and minimum inhibitory concentration (MIC) against various bacteria  Singh et al. (2023).
@@ 3 Continual learning scenarios
Depending upon what kind of change in the sequence of data is expected, the existing CL scenarios can be broadly categorized into five categories, viz., (a) data-incremental, (b) class-incremental, (c) task-incremental, (d) domain-incremental, and (e) hybrid. Figure 4 shows the percentages of the above-mentioned incremental CL scenarios exploited in recent years. The statistics reflect that domain-incremental is highly explored with a 51.9% ratio out of five major settings, followed by the class-incremental settings having 26.9% works and data-incremental being the least challenging and rarely explored for different medical image analysis applications.
Table 2: List of various instance incremental scenarios in literature
@@
Ravishankar et al. (2019) Pneumothorax identification from X-ray images (2 Ep.) 4 batches each of 2K were created from a subset of ChestXRay (Wang et al. 2017)
Kaustaban et al. (2022) Histopathology data-based tumor classification (4 Ep.) Ep. was created in instance incremental fashion from colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes, (i.e., all 9 classes are present in each episode)
3.1 Instance incremental scenarios
Instance Incremental Scenario (IIS), also termed as data incremental scenario, involves the continuous learning process where the model keeps adapting to incoming data streams that come from the same data distribution. This scenario doesn’t involve dealing with entirely new categories or significant changes in data patterns, making it generally the least challenging among all other CL scenarios. In the medical domain, especially digital pathology, a common practice is for expert pathologists to annotate a dataset in multiple stages or batches Kaustaban et al. (2022). Each of these annotation batches can be thought of as representing samples from the same underlying data distribution, ensuring some consistency in the data source across different batches. Thus, IIS scenarios deal with the condition where the samples of a dataset are sequentially made available. Owning to its simplicity, this category is not very popular. Only a few classification applications are explored in this scenario, as mentioned in Table 2. For example, in the breast cancer classification application, Chen and Tang (2022) set up IIS on the BreakHis Spanhol et al. (2015) dataset. More and more samples from each of the malignant and benign classes are added over the subsequent episodes. Ravishankar et al. (2019) try to exhibit a scenario for pneumothorax detection in a hospital where data arrives in an incremental fashion. They sequentially fed 2000 samples in 4 subsequent episodes from an X-ray dataset for Pneumothorax identification consisting of a total of 8000 samples.
@@ 3.1 Instance incremental scenarios
Ozdemir et al. (2018) MRI based Humerus and scapula segmentation (2 Ep.) Incrementally adding anatomical structure, private data
@@
Ozdemir and Goksel (2019) MRI based knee segmentation (2 Ep.) SKI10 MICCAI Grand Challenge (Heimann et al. 2010)
@@
Li et al. (2020) Dermoscopic images based skin disease classification (4-20 Ep.) (a) CIS for Skin8 dataset having total 8 classes on which 2 classes per episode used, (b) CIS for Skin40 dataset from ISIC2019 (Tschandl et al. 2018) consists of 40 classes of skin disease images collected from the internet, here 2/5/10 classes per episode used in 3 exp, (c) non-medical dataset CIFAR100 with 100 classes, here 5/10/20 classes per episode used in 3 exp
@@
Liu et al. (2022) Class incremental segmentation of abdomen organs (4 Ep.) 4 Ep. created for incremental segmentation of liver, spleen, pancreas, right kidney, and left kidney organs from CT datasets (Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
@@
Derakhshani et al. (2022) CIS: disease classification (4 Ep.) 4 Ep. created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (taken from MedMNIST repository (Yang et al. 2023a))
@@
Akundi and Sivaswamy (2022) Chest X-ray classification in CIS (5 Ep.) 5 classes from CheXpert dataset
@@
Kaustaban et al. (2022) Histopathology data based tumor classification (4 Ep.) 4 Ep. were created in CIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total 9 classes
@@
Chen and Tang (2022) Histopathology data-based breast cancer classification (4 Ep.) 4 Ep. created in CIS from BreakHis dataset
@@
Zhang et al. (2023b) (a) CT based abdomen, gastrointestinal, and other organ segmentation (7, 3, 4 classes in 3 Ep.), (b) abdomen segmentation to liver tumor segmentation task (13 classes, 1 class in 2 Ep.) (a) Abdomen segmentation from JHH (Xia et al. 2022) (private) (b) 13 class abdomen segmentation dataset as BTCV (Landman et al. 2015) and liver tumor segmentation data as LiTS (Bilic et al. 2023)
@@
Ji et al. (2023) 3D CT scan based whole-body organs segmentation (4 Ep. in 2 permutations of episodes) TotalSegmentator (Wasserthal et al. 2023) (103 classes: various organs), 3 private datasets: ChestOrgan (31 classes: chest scans), HNOrgan (13 classes: head and neck scans), EsoOrgan (1 class: esophageal cancer)
@@
Chee et al. (2023) Cancer classification (3,5 Ep.), diabetic retinopathy classification (2 Ep.), skin lesions classification (3,4 Ep.) (a) Histopathology data based human colorectal cancer classification: CCH5000 (Kather et al. 2016) (total 8 classes, Exp=(4,1,1,1,1), Exp=(4,2,2)), (b) diabetic retinopathy (DR) classification using retinal images: EyePACS (kag ) (total 5 classes: no DR, mild DR, moderate DR, severe DR and proliferative DR, Exp=(3,2)), (c) pigmented skin lesions classification: HAM10000 (Tschandl et al. 2018) (total 7 classes: Exp=(4,1,1,1), Exp=(4,2,1))
@@
Zhang et al. (2023c) Disease classification in CIS (4-10 Ep.) (a) skin lesions classification with Skin8 dataset (International Skin Imaging Collaboration (ISIC) (Tschandl et al. 2018)) having 8 classes distributed over 4 Ep., (b) Path16, a pathology image collection from various public sources: total 16 classes distributed over 7 episodes, (c) CIFAR100: non-medical images, 100 classes distributed over 5-10 episodes
(Bai et al. 2023) Surgical visual-question localized-answering EndoVis18, EndoVis17, M2CAI
@@
Sadafi et al. (2023) Microscopic images based WBC classification (4 Ep.) With 3 different datasets, 3 separate class-incremental scenarios each having 4 episodes: (a) (3,3,3,4) classes from Matek-19 (Matek et al. 2019), (b) (3,3,3,4) classes from INT-20, and (c) (3,2,2,3) classes from Acevedo-20 (Acevedo et al. 2020)
@@ 3.2 Class incremental scenarios
Class Incremental Scenario (CIS) refers to a situation where the aim is to adapt the model to accommodate novel classes from new data streams. Given that each batch of these unseen classes offers knowledge that largely differs from the previous batches, this scenario is anticipated to be the most challenging among continual learning scenarios. An instance of this challenge is incorporating new types of tissues into a tumor classification model or introducing novel cell types into a model designed for AI-assisted diagnosis in cell detection. In this scenario, there is only one incremental task that gets additions of new classes in subsequent episodes. Some works assume mutually exclusive classes, whereas some overlapping classes.
Multi-class datasets with more than 5-6 classes are mainly considered to create a class incremental scenario. There can be an equal or variable number of classes across the episodes. Chen and Tang (2022) propose a breast cancer classification model where they iteratively include malignant and benign classes with one sub-class from each category. This scenario is very popular for segmentation applications where the model may be demanded to segment more and more classes over time. An exhaustive list of class incremental scenarios explored in literature is provided via Table 3. References, along with detailed segmentation or classification applications, are tabulated.
@@
Ravishankar et al. (2019) Chest x-ray view classification (2 Ep.) Shift from task (4ch vs. PLAX) to task (2ch vs. PSAX)
Kaustaban et al. (2022) Histopathology data based tumor classification (4 Ep.) 4 Ep. were created in TIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes
Derakhshani et al. (2022) Disease classification (4 Ep.) 4 Ep. created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (MedMNIST repository (Yang et al. 2023a))
Bera et al. (2023) Three tasks involving different organs for three different binary segmentation tasks (3 Ep.) MRI data for binary prostate, spleen, hippocampus segmentation: Promise12 (Litjens et al. 2014) (prostate) $\rightarrow$ MSD (NCI ) (spleen) $\rightarrow$ Drayd (Denovellis et al. 2021) (hippocampus)